# The hereditariness problem for the Černý conjecture

### v1

@@ 6 The radical ideal and the radical congruence
This section is mainly devoted to the non-semisimple case, with special attention to the radical ideal $\text{Rad}(\mathcal{A}^{*})$, its computation, and the associated congruence, studied via the Galois correspondence introduced in the previous section. In particular, we establish a connection between the index of nilpotency of $\text{Rad}(\mathcal{A}^{*})$ and the height of the lattice $\text{Cong}(\mathcal{A})$.
We start with the following remark: given $I\subseteq\mathcal{A}^{*}$ an ideal, if $\theta:\text{M}(\mathcal{A})\rightarrow\mathcal{A}^{*}$ is the Rees morphism we can define:
$\rho(I):=\rho(\theta^{-1}(I))=\rho(I\cup\text{Syn}(\mathcal{A}))=\bigcap_{u\in I\cup\text{Syn}(\mathcal{A})}\text{ker}(u)=\bigcap_{u\in I}\text{ker}(u)$
which allows us to restrict the definition of $\operatorname{Id}(\mathcal{A})$ to ideals of $\mathcal{A}^{*}$. From now on, by the above expression we will not make any difference between $\rho(I)$ and $\rho(\theta^{-1}(I))$ for a given ideal $I\subseteq\mathcal{A}^{*}$. We begin with the study of the congruence given by the radical ideal, denoted by $\rho:=\rho(\text{Rad}(\mathcal{A}))$. We will refer to such congruence as radical congruence. Throghout the rest of this work, by abuse of notation we will use $\text{Rad}(\mathcal{A})$ also to refer to $\text{Rad}(\mathcal{A}^{*})$ and for any $I\subseteq\Sigma^{*}$ two-sided ideal, we define $\rho(I):=\rho(\pi(I))$ where $\pi:\Sigma^{*}\rightarrow\text{M}(\mathcal{A})$ is the usual epimorphism.
@@
Let $\mathcal{A}$ be a DFA. Then, either $\mathcal{A}$ semisimple or $\rho$ is nontrivial.
@@
Assume that $\mathcal{A}$ is not semisimple and so let $u\in\text{Rad}(\mathcal{A})\setminus\text{Syn}(\mathcal{A})$. Then, there exist distinct states $p,q\in Q$ such that $p\cdot u\neq q\cdot u$, hence $(p,q)\notin\rho$ which implies that $\rho\neq\Delta_{\mathcal{A}}$.
Let now $m\in\mathbb{N}$ be the nihilpotency index of $\text{Rad}(\mathcal{A})/\text{Syn}(\mathcal{A})$ in $\Sigma^{*}/\text{Syn}(\mathcal{A})$. By definition, we may find $u:=u_{1}\ldots u_{m-1}\in\text{Rad}(\mathcal{A})\setminus\text{Syn}(\mathcal{A})$ such that $u_{i}\in\text{Rad}(\mathcal{A})$ for every $i\in\{1,\ldots,m-1\}$, and $uv\in\text{Syn}(\mathcal{A})$ for every $v\in\text{Rad}(\mathcal{A})$. Since $u\notin\text{Syn}(\mathcal{A})$, we have that $Q\cdot u\supseteq\{p,q\}$ with $p\neq q$, and thus $p\cdot v=q\cdot v$ for every $v\in\text{Rad}(\mathcal{A})$. Therefore, $(p,q)\in\rho$, and so we conclude that $\rho\neq\Delta_{\mathcal{A}}$. ∎
@@
Recall that the height of a (finite) lattice (or poset) is the maximum length of its chains. The following theorem connects the height of the congruence lattice with the radical of the automaton, in particular by showing that the height bounds the index of nilpotency of $\text{Rad}(\mathcal{A})$.
@@
Let $\mathcal{A}$ be a synchronizing DFA, and let $m$ be the height of $\text{Cong}(\mathcal{A})$. Then
@@
In other words, the nilpotency index of $\text{Rad}(\mathcal{A})$ is bounded above by the height of the congruence lattice minus one, and this bound is tight.
@@
Let $\rho_{0}:=\Delta_{\mathcal{A}},\ \mathcal{A}_{0}:=\mathcal{A}$ and $\forall i\geq 0$ define the following:
$\rho_{i+1}:=\rho_{\mathcal{A}_{i}}=\bigcap_{u\in\text{Rad}(\mathcal{A}_{i})}\text{ker}_{\mathcal{A}_{i}}(u),\text{ where }\mathcal{A}_{i}:=\mathcal{A}_{i-1}/\rho_{i}$
Observe that the lifting congruence $\overline{\rho}_{i}$ belongs to $\text{Cong}(\mathcal{A})$ for every $i$, as defined in Proposition 4.6. Since the automaton is finite, there exists an index $k$ such that $\mathcal{A}_{k}$ is semisimple. Moreover, because a maximal non-trivial congruence $\sigma\in\text{Cong}(\mathcal{A})$ yields a simple quotient $\mathcal{A}/\sigma$, it follows that for some $k$, either $\overline{\rho}_{k}$ is maximal or $\mathcal{A}_{k-1}$ is semisimple. In the first case we still have that $\mathcal{A}_{k}$ is semisimple.
Let $k$ be such integer and observe that by Lemma 5.4 we have:
$\begin{split}\text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})&\subseteq\text{Syn}(\mathcal{A}_{k-1})\\ \text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})\text{Rad}(\mathcal{A}_{k-2})&\subseteq\text{Syn}(\mathcal{A}_{k-2})\\ &\dots\\ \text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})\dots\text{Rad}(\mathcal{A})&\subseteq\text{Syn}(\mathcal{A})\\ \text{Rad}(\mathcal{A})^{k+1}&\subseteq\text{Syn}(\mathcal{A})\end{split}$
thus we have a bound on the nihilpotency index of $\text{Rad}(\mathcal{A})$ in $\Sigma^{*}/\text{Syn}(\mathcal{A})$. By maximality of $m$, together with the fact that the $\rho_{1},\ldots,\rho_{k}$ are non trivial, we have $k+2\leq m$, and therefore
@@
which concludes the proof.
@@
It is easy to check that $\rho_{1}=\ker(a)$ and $\rho_{2}=\ker(a^{2})$, with $\mathcal{S}_{3}/\rho_{2}$ being a simple automaton. It is also easy to check that the chain $\Delta_{\mathcal{S}_{3}}\subseteq\rho_{1}\subseteq\rho_{2}\subseteq\nabla_{\mathcal{S}_{3}}$ realizes the full poset $\text{Cong}(\mathcal{A})$, from which we have m1=3m-1=3 and $\text{Rad}(\mathcal{S}_{3})^{3}=\text{Syn}(\mathcal{S}_{3})$ with $\text{Rad}(\mathcal{S}_{3})^{2}\neq\text{Syn}(\mathcal{S}_{3})$.
@@
$\rho_{1}\subseteq\ldots\subseteq\rho_{n-1}\ \text{ with }\ \rho_{i}=\ker(a^{i})$
@@
In the setting of the previous theorem, observe that since $\mathcal{A}_{k}$ is semisimple, one might hope to apply this construction to prove Conjecture 4.15 by induction along the chain of congruences $\rho_{i}$, attempting to lift a Černý reset word from $\mathcal{A}_{i+1}$ to $\mathcal{A}_{i}$. Fix $i\in\{1,\ldots,k-1\}$ and assume that $\rho_{i+1}$ admits a $1$-, 22-, or $3$-class as a congruence of $\mathcal{A}/\rho_{i}$. By Proposition 3.1, we can easily construct a synchronizing word for $\mathcal{A}_{i}$ whenever one for $\mathcal{A}_{i+1}$ is given. Hence, in this case we would be able to solve the hereditariness problem Assume instead that, for every $i\in\{0,\ldots,k-1\}$, the congruence $\rho_{i+1}$ admits only classes of cardinality greater than $4$. In this situation, we were not able to solve the problem using any combinatorial technique to lift a Černý reset word. Nevertheless, one may observe that $|\mathcal{A}_{i}|\leq|\mathcal{A}_{i-1}|/4$ for every $i\in\{1,\ldots,k\}$, and thus $k+1\leq\log_{4}(n)$, showing that the nilpotency index of $\text{Rad}(\mathcal{A})$ is bounded by $\log_{4}(n)$.
@@
and thus $\text{Rad}(\mathcal{A})\subseteq\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)$. Observe now that, being $\text{Rad}(\mathcal{A}/\sigma)$ a nihilpotent ideal of $(\mathcal{A}/\sigma)^{*}$, there exists $k\geq 1$ such that $\text{Rad}(\mathcal{A}/\sigma)^{k}\subseteq\text{Syn}(\mathcal{A}/\sigma)$ and thus:
@@
and thus $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)$ is a nihilpotent ideal of $\mathcal{A}^{*}$. By the fact that $\text{Rad}(\mathcal{A})$ is the biggest nihilpotent ideal of $\mathcal{A}^{*}$, we have $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)\subseteq\text{Rad}(\mathcal{A})$ and thus $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)=\text{Rad}(\mathcal{A})$. The last statement follows immediately by $\rho\in\text{Cong}_{\rho}(\mathcal{A})$. ∎
@@ 7.1 The non-semisimple case
where $\rho=\rho(\text{Rad}(\mathcal{A}))$, and where $\text{Cong}(\mathcal{A})\setminus\{\Delta_{\mathcal{A}}\}$ admits a minimum.
Recall that associated with $\text{Rad}(\mathcal{A})$ there is the radical congruence $\rho=\rho(\text{Rad}(\mathcal{A}))$. Moreover, since $\mathcal{A}$ is radical, $\text{Cong}(\mathcal{A})\setminus\{\Delta_{\mathcal{A}}\}$ admits a minimum, which we denote by $\sigma$. In this setting we obtain, by Proposition 6.5, that
@@
one may check that $ba^{2}b$ is a non-reset radical word, and that the only non-trivial congruence $\sigma$ is generated by the following partition $\big\{\{q_{0},q_{1}\},\{q_{2}\},\{q_{3}\}\big\}$, with $\mathcal{A}/\sigma\cong\mathcal{C}_{3}$, the Černý automaton of order $3$. Thus, $\mathcal{A}$ is not semisimple and $\rho(\text{Rad}(\mathcal{A}))=\sigma$.
Now let $u\in\text{Syn}(\mathcal{A}/\sigma)\setminus\text{Syn}(\mathcal{A})$. We must then have $Q\cdot u=\{q_{0},q_{1}\}$. Observe that for any $v\in\Sigma^{*}$,
@@
But in this case $v\notin\text{Syn}(\mathcal{A}/\sigma)$. Hence we must have $q_{0}\cdot u=q_{1}\cdot u$, and therefore $u\in I(\sigma)$. This shows that $\text{Syn}(\mathcal{A}/\sigma)\subseteq I(\sigma)$, and from Proposition 6.5 we obtain
$\text{Rad}(\mathcal{A})\;=\;I(\sigma)\cap\text{Syn}(\mathcal{A}/\sigma)\;=\;\text{Syn}(\mathcal{A}/\sigma)\;=\;\text{Rad}(\mathcal{A}/\rho).$
@@
  • Assume that $\text{Cong}(\mathcal{A})$ has a non-trivial minimum. Then by hypothesis, since $\mathcal{A}$ is not radical, we must have that $\text{Rad}(\mathcal{A}/\rho)=\text{Rad}(\mathcal{A})$. By the induction hypothesis, let $u\in\text{Syn}(\mathcal{A}/\rho)$ be a word satisfying the Černý conjecture. If $\rho$ admits a 11- or 22-class, we are done by means of Proposition 3.1 by strongly connectedness. Otherwise, observe that:
@@
  • Assume that $\text{Cong}(\mathcal{A})$ admits two non-trivial minimal congruences $\sigma_{1},\sigma_{2}$. If one of them admits a $1$-class (say $\sigma_{1}$, without loss of generality), the claim follows by applying the induction hypothesis to $\mathcal{A}/\sigma_{1}$ and then linearly synchronizing the resulting word. Otherwise, by Proposition 4.10 we have
@@
We conclude this subsection by presenting a partial structural result concerning the 0-minimal ideals of $\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ in the radical case. Our motivation for focusing on these ideals stems from their connection with the Wedderburn–Artin decomposition of the automaton, as highlighted in Remark 2.4. We need first the following lemma. Let $\theta:\mathcal{A}^{*}\to\mathcal{A}^{*}/\text{Rad}(\mathcal{A})\cong\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ denote the Rees morphism. For an ideal $I\subseteq\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$, we define
$\rho(I):=\rho\!\bigl(\theta^{-1}(I)\bigr).$
@@
Observe that $I_{1}I_{2}=0$ can be rephrased by abuse of notation as $I_{1}I_{2}\subseteq\text{Rad}(\mathcal{A})$ as ideals of $\mathcal{A}^{*}$. Then, by another abuse of notation we can rephrase the latter inclusion as:
@@
Let $\mathcal{A}$ be a radical DFA, $\sigma\in\text{Cong}(\mathcal{A})$ the non-trivial minimum congruence. Then $\text{Syn}(\mathcal{A}/\sigma)$ contains exactly one 0-minimal ideal seen as an ideal of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$.
@@
First, observe that $\text{Syn}(\mathcal{A}/\sigma)/\text{Rad}(\mathcal{A})$ has to be non-trivial, for if otherwise we would have that $\text{Syn}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A})$ as ideals in $\text{M}(\mathcal{A})$ and by Lemma 4.9 this implies $\text{Rad}(\mathcal{A})=\text{Rad}(\mathcal{A}/\sigma)$, against the hypothesis of $\mathcal{A}$ being radical. Let $I_{1},I_{2}\subseteq\text{Syn}(\mathcal{A}/\sigma)/\text{Rad}(\mathcal{A})$ be two non trivial minimal ideals. By means of the previous lemma we have that either $\rho(I_{1})$ or $\rho(I_{2})$ has to be non-trivial, without loss of generality assume $\rho(I_{1})$ non-trivial. Then, by minimality we have $\sigma\subseteq\rho(I_{1})$, from which we deduce

### v2

@@ 6 The radical ideal and the radical congruence
This section is mainly devoted to the non-semisimple case, with special attention to the radical ideal $\text{Rad}(\mathcal{A}^{\star})$, its computation, and the associated congruence, studied via the Galois correspondence introduced in the previous section. In particular, we establish a connection between the index of nilpotency of $\text{Rad}(\mathcal{A}^{\star})$ and the height of the lattice $\text{Cong}(\mathcal{A})$.
We start with the following remark: given $I\subseteq\mathcal{A}^{\star}$ a two-sided ideal, if $I=0$ we define $\rho(I):=\nabla_{\mathcal{A}}$. Otherwise, if $I\neq 0$ we define $\rho(I):=\rho(\theta^{-1}(I))$ where $\theta:\text{M}(\mathcal{A})\rightarrow\mathcal{A}^{\star}$ is the Rees morphism. Note that:
$\rho(I)=\rho(\theta^{-1}(I))=\rho((I\setminus\{0\})\cup\text{Syn}(\mathcal{A}))=\bigcap_{u\in(I\setminus\{0\})\cup\text{Syn}(\mathcal{A})}\text{ker}(u)=\bigcap_{u\in I\setminus\{0\}}\text{ker}(u)$
which allows us to restrict the definition of $\operatorname{Id}(\mathcal{A})$ to ideals of $\mathcal{A}^{\star}$. From now on, by the above expression we will not make any difference between $\rho(I)$ and $\rho(\theta^{-1}(I))$ for a given ideal $I\subseteq\mathcal{A}^{\star}$. We begin with the study of the congruence given by the radical ideal, denoted by $\rho:=\rho(\text{Rad}(\mathcal{A}))$. We will refer to such congruence as radical congruence. Throghout the rest of this work, by abuse of notation we will use $\text{Rad}(\mathcal{A})$ also to refer to $\text{Rad}(\mathcal{A}^{\star})$ and for any $I\subseteq\Sigma^{*}$ two-sided ideal, we define $\rho(I):=\rho(\pi(I))$ where $\pi:\Sigma^{*}\rightarrow\text{M}(\mathcal{A})$ is the usual epimorphism.
@@
Let $\mathcal{A}$ be a DFA. Then, either $\mathcal{A}$ semisimple, in which case $\rho=\nabla_{\mathcal{A}}$, or $\rho\notin\{\nabla_{\mathcal{A}},\Delta_{\mathcal{A}}\}$.
@@
If $\mathcal{A}$ is semisimple we clearly have that $\rho=\rho(\text{Rad}(\mathcal{A}))=\rho(\text{Syn}(\mathcal{A}))=\nabla_{\mathcal{A}}$. Assume that $\mathcal{A}$ is not semisimple and so let $u\in\text{Rad}(\mathcal{A})\setminus\text{Syn}(\mathcal{A})$. Then, there exist distinct states $p,q\in Q$ such that $p\cdot u\neq q\cdot u$, hence $(p,q)\notin\rho$ which implies that $\rho\neq\nabla_{\mathcal{A}}$.
Let now $m\in\mathbb{N}$ be the nilpotency index of $\text{Rad}(\mathcal{A}^{\star})$. By definition, we may find $u:=u_{1}\ldots u_{m-1}\in\text{Rad}(\mathcal{A})\setminus\text{Syn}(\mathcal{A})$ such that $u_{i}\in\text{Rad}(\mathcal{A})$ for every $i\in\{1,\ldots,m-1\}$, and $uv\in\text{Syn}(\mathcal{A})$ for every $v\in\text{Rad}(\mathcal{A})$. Since $u\notin\text{Syn}(\mathcal{A})$, we have that $Q\cdot u\supseteq\{p,q\}$ with $p\neq q$, and thus $p\cdot v=q\cdot v$ for every $v\in\text{Rad}(\mathcal{A})$. Therefore, $(p,q)\in\rho$, and so we conclude that $\rho\neq\Delta_{\mathcal{A}}$. ∎
@@
Recall that the height of a (finite) lattice (or poset) is the maximum length of its chains. We will now prove a theorem connecting the height of the congruence lattice with the radical of the automaton, in particular by showing that the height bounds the index of nilpotency of $\text{Rad}(\mathcal{A})$. We first need to state the following lemma:
Lemma 6.3.
Let $\mathcal{A}$ be a non-semisimple automaton, $\rho=\rho(\text{Rad}(\mathcal{A}))$ as above. We have that:
$\text{Syn}(\mathcal{A}/\rho)\cdot\text{Rad}(\mathcal{A})\subseteq\text{Syn}(\mathcal{A}).$
Proof.
It is enough to observe that, being $\rho:\operatorname{Id}(\mathcal{A})\rightarrow\text{Cong}(\mathcal{A})$ an antimorphism of a Galois connection, we have that $\text{Rad}(\mathcal{A})\subseteq I(\rho(\text{Rad}(\mathcal{A})))=I(\rho)$, and thus by Proposition 5.9 we obtain:
$\text{Syn}(\mathcal{A}/\rho)\cdot I(\rho)\subseteq\text{Syn}(\mathcal{A})\Rightarrow\text{Syn}(\mathcal{A}/\rho)\cdot\text{Rad}(\mathcal{A})\subseteq\text{Syn}(\mathcal{A})$
which concludes. ∎
@@
Let $\mathcal{A}$ be a non-semisimple synchronizing DFA, and let $m$ be the height of $\text{Cong}(\mathcal{A})$. Then
@@
In other words, the nilpotency index of $\text{Rad}(\mathcal{A})$ is bounded above by the height of the congruence lattice minus one. This bound is tight.
@@
Let $\rho_{0}:=\Delta_{\mathcal{A}}$ and for each $i\geq 0$ define iteratively the following congruences:
$\rho_{i+1}:=\rho_{\mathcal{A}_{i}}=\bigcap_{u\in\text{Rad}(\mathcal{A}_{i})}\text{ker}_{\mathcal{A}_{i}}(u),\ \text{ where }\ \mathcal{A}_{0}:=\mathcal{A}\ \text{ and }\ \mathcal{A}_{i}:=\mathcal{A}_{i-1}/\rho_{i}\ \text{ for }\ i\geq 1$
Observe that the lifting congruence $\overline{\rho}_{i}$ belongs to $\text{Cong}(\mathcal{A})$ for every $i$, as defined in Proposition 4.3. Furthermore, by iterating the same proposition we conclude that $\overline{\rho}_{i}\subseteq\overline{\rho}_{i+1}$ for every $i\geq 0$. Since $\text{Cong}(\mathcal{A})$ is finite, there exists an index $k$ such that $\mathcal{A}_{k}$ is semisimple. Indeed, assume by contradiction that $\mathcal{A}_{i}$ is not semisimple for every $i\geq 0$. By Lemma 6.1, we have that $\rho_{i+1}$ is a non-trivial congruence of $\mathcal{A}_{i}$, and thus $\overline{\rho}_{i}\subsetneq\overline{\rho}_{i+1}$ for every $i\geq 0$. This gives us an infinite chain in $\text{Cong}(\mathcal{A})$:
$\overline{\rho}_{0}\subsetneq\overline{\rho}_{1}\subsetneq\overline{\rho}_{2}\subsetneq\ldots\subsetneq\overline{\rho}_{i}\subsetneq\ldots$
contradicting the finiteness of $\text{Cong}(\mathcal{A})$. Let $k$ be such integer and observe that we have a chain in $\text{Cong}(\mathcal{A})$ as follows
$\overline{\rho}_{0}=\Delta_{\mathcal{A}}\subsetneq\overline{\rho}_{1}\subsetneq\ldots\subsetneq\overline{\rho}_{k}\subsetneq\overline{\rho}_{k+1}=\nabla_{\mathcal{A}}$
where again $\overline{\rho}_{k+1}=\nabla_{\mathcal{A}}$ is a consequence of Lemma 6.1. By Lemma 6.3 we have:
$\text{Syn}(\mathcal{A}_{i})\text{Rad}(\mathcal{A}_{i-1})\subseteq\text{Syn}(\mathcal{A}_{i-1})$
and by applying the above inclusion iteratively for any $i$ we obtain:
$\begin{split}\text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})&\subseteq\text{Syn}(\mathcal{A}_{k-1})\\ \text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})\text{Rad}(\mathcal{A}_{k-2})&\subseteq\text{Syn}(\mathcal{A}_{k-2})\\ &\dots\\ \text{Syn}(\mathcal{A}_{k})\text{Rad}(\mathcal{A}_{k-1})\dots\text{Rad}(\mathcal{A}_{0})&\subseteq\text{Syn}(\mathcal{A})\\ \text{Rad}(\mathcal{A})^{k+1}&\subseteq\text{Syn}(\mathcal{A})\end{split}$
where the last line is obtained by observing that $\text{Rad}(\mathcal{A})\subseteq\text{Rad}(\mathcal{A}_{i})$ for every $i$ and $\text{Syn}(\mathcal{A}_{k})=\text{Rad}(A_{k})$. Thus we have a bound on the nilpotency index of $\text{Rad}(\mathcal{A}^{\star})$. By maximality of $m$, together with the fact that the $\overline{\rho}_{1},\ldots,\overline{\rho}_{k}$ are non-trivial, we have $k+2\leq m$, and therefore
@@
which concludes the first part of the proof.
@@
It is easy to check that $\overline{\rho}_{1}=\ker(a)$ and $\overline{\rho}_{2}=\ker(a^{2})$, with $\mathcal{S}_{4}/\rho_{2}$ being a simple automaton. It is also easy to check that the chain $\Delta_{\mathcal{S}_{4}}\subseteq\overline{\rho}_{1}\subseteq\overline{\rho}_{2}\subseteq\nabla_{\mathcal{S}_{4}}$ realizes the full poset $\text{Cong}(\mathcal{A})$, from which we have m1=3m-1=3 and $\text{Rad}(\mathcal{S}_{4})^{3}=\text{Syn}(\mathcal{S}_{4})$ with $\text{Rad}(\mathcal{S}_{4})^{2}\neq\text{Syn}(\mathcal{S}_{4})$.
@@
$\overline{\rho}_{1}\subseteq\ldots\subseteq\overline{\rho}_{n-1}\ \text{ with }\ \overline{\rho}_{i}=\ker(a^{i})$
@@
In the setting of the previous theorem, observe that since $\mathcal{A}_{k}$ is semisimple, one might hope to apply this construction to prove Conjecture 4.11 by induction along the chain of congruences $\rho_{i}$, attempting to lift a Černý reset word from $\mathcal{A}_{i+1}$ to $\mathcal{A}_{i}$. Fix $i\in\{1,\ldots,k-1\}$ and assume that $\rho_{i+1}$ admits a $1$-, 22-, or $3$-class as a congruence of $\mathcal{A}/\rho_{i}$. By Proposition 3.1, we can easily construct a synchronizing word for $\mathcal{A}_{i}$ whenever one for $\mathcal{A}_{i+1}$ is given. Hence, in this case we would be able to solve the hereditariness problem. Assume instead that, for every $i\in\{0,\ldots,k-1\}$, the congruence $\rho_{i+1}$ admits only classes of cardinality greater than $4$. In this situation, we were not able to solve the problem using any combinatorial technique to lift a Černý reset word. Nevertheless, one may observe that $|\mathcal{A}_{i}|\leq|\mathcal{A}_{i-1}|/4$ for every $i\in\{1,\ldots,k\}$, and thus $k+1\leq\log_{4}(n)$, showing that the nilpotency index of $\text{Rad}(\mathcal{A})$ is bounded by $\log_{4}(n)$.
@@
and thus $\text{Rad}(\mathcal{A})\subseteq\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)$. Observe now that, being $\text{Rad}(\mathcal{A}/\sigma)$ a nilpotent ideal of $(\mathcal{A}/\sigma)^{\star}$, there exists $k\geq 1$ such that $\text{Rad}(\mathcal{A}/\sigma)^{k}\subseteq\text{Syn}(\mathcal{A}/\sigma)$ and thus:
@@
where $\text{Syn}(\mathcal{A}/\sigma)\cdot I(\sigma)\subseteq\text{Syn}(\mathcal{A})$ is due to Lemma 5.4. This gives us that $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)$ is a nilpotent ideal of $\mathcal{A}^{\star}$. By the fact that $\text{Rad}(\mathcal{A})$ is the biggest nilpotent ideal of $\mathcal{A}^{\star}$, we have $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)\subseteq\text{Rad}(\mathcal{A})$ and thus $\text{Rad}(\mathcal{A}/\sigma)\cap I(\sigma)=\text{Rad}(\mathcal{A})$. The last statement follows immediately by $\rho\in\text{Cong}_{\rho}(\mathcal{A})$. ∎
@@ 7.1 The non-semisimple case
where $\rho=\rho(\text{Rad}(\mathcal{A}))$, and where $\mathcal{A}$ admits a monolith (in other words, $\text{Cong}(\mathcal{A})\setminus\{\Delta_{\mathcal{A}}\}$ has a minimum).
Recall that associated with $\text{Rad}(\mathcal{A})$ there is the radical congruence $\rho=\rho(\text{Rad}(\mathcal{A}))$. Moreover, since $\mathcal{A}$ is radical, it admits a monolith $\sigma$. In this setting we obtain, by Proposition 6.6, that
@@
one may check that $ba^{2}b$ is a non-reset radical word, and that the only non-trivial congruence $\sigma$ is generated by the following partition $\big\{\{q_{0},q_{1}\},\{q_{2}\},\{q_{3}\}\big\}$, with $\mathcal{A}/\sigma\cong\mathcal{C}_{3}$, the Černý automaton of order $3$. Thus, $\mathcal{A}$ is not semisimple and $\rho(\text{Rad}(\mathcal{A}))=\sigma$. Let us now show that $\mathcal{A}$ is not radical, by showing that $\text{Rad}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A})$. Now let $u\in\text{Syn}(\mathcal{A}/\sigma)\setminus\text{Syn}(\mathcal{A})$. We must then have $Q\cdot u=\{q_{0},q_{1}\}$. Observe that for any $v\in\Sigma^{*}$,
@@
But in this case $v\notin\text{Syn}(\mathcal{A}/\sigma)$. Therefore, we must have $q_{0}\cdot u=q_{1}\cdot u$, which implies $u\in I(\sigma)$. This shows that $\text{Syn}(\mathcal{A}/\sigma)\subseteq I(\sigma)$, and from Proposition 6.6 we obtain
$\text{Rad}(\mathcal{A})\;=\;I(\sigma)\cap\text{Syn}(\mathcal{A}/\sigma)\;=\;\text{Syn}(\mathcal{A}/\sigma)\;=\;\text{Rad}(\mathcal{A}/\sigma).$
@@
  • Assume that $\mathcal{A}$ has a monolith. Then by hypothesis, since $\mathcal{A}$ is not radical, we must have that $\text{Rad}(\mathcal{A}/\rho)=\text{Rad}(\mathcal{A})$. By the induction hypothesis, let $u\in\text{Syn}(\mathcal{A}/\rho)$ be a Černý reset word. If $\rho$ admits a 1- or 2-class, we are done by means of Proposition 3.1 by strongly connectedness. Otherwise, by Lemma 6.3 we have that:
@@
  • Assume that $\text{Cong}(\mathcal{A})$ admits two non-trivial minimal congruences $\sigma_{1},\sigma_{2}$. If one of them admits a $1$-class (say $\sigma_{1}$, without loss of generality), the claim follows by Proposition 3.1. Otherwise, by Proposition 4.6 we have
@@
We conclude this subsection by presenting a partial structural result concerning the 0-minimal ideals of $\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ in the radical case. Our motivation for focusing on these ideals stems from their connection with the Wedderburn–Artin decomposition of the automaton, as highlighted in Remark 2.4. We need first the following lemma. Let $\xi:\mathcal{A}^{\star}\to\mathcal{A}^{\star}/\text{Rad}(\mathcal{A})\cong\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$ denote the Rees morphism. For an ideal $I\subseteq\mathrm{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$, we define
$\rho(I):=\rho\!\bigl(\xi^{-1}(I)\bigr).$
@@
Observe that $I_{1}I_{2}=0$ can be rephrased by abuse of notation as $I_{1}I_{2}\subseteq\text{Rad}(\mathcal{A})$ as ideals of $\mathcal{A}^{\star}$. Then, by another abuse of notation we can rephrase the latter inclusion as:
@@
Let $\mathcal{A}$ be a radical DFA, $\sigma\in\text{Cong}(\mathcal{A})$ its monolith. Then $\text{Syn}(\mathcal{A}/\sigma)$ contains exactly one 0-minimal ideal seen as an ideal of $\text{M}(\mathcal{A})/\text{Rad}(\mathcal{A})$.
@@
First, observe that $\text{Syn}(\mathcal{A}/\sigma)/\text{Rad}(\mathcal{A})$ has to be non-trivial, for if otherwise we would have that $\text{Syn}(\mathcal{A}/\sigma)\subseteq\text{Rad}(\mathcal{A})$ as ideals in $\text{M}(\mathcal{A})$ and by Lemma 4.5 this implies $\text{Rad}(\mathcal{A})=\text{Rad}(\mathcal{A}/\sigma)$, against the hypothesis of $\mathcal{A}$ being radical. Let $I_{1},I_{2}\subseteq\text{Syn}(\mathcal{A}/\sigma)/\text{Rad}(\mathcal{A})$ be two non-trivial minimal ideals. By means of the previous lemma we have that either $\rho(I_{1})$ or $\rho(I_{2})$ has to be non-trivial, without loss of generality assume $\rho(I_{1})$ non-trivial. Then, by minimality we have $\sigma\subseteq\rho(I_{1})$, from which we deduce